Over the weekend, Bob and Daphen gave the PRM a 48hr. vacuum bake. This afternoon Suresh and I placed it back in the wire in the tower. We used the microscope-on-translation-stage technique to make sure the scribe lines on each side of the optic are at the same height, and then secured the PRM using all of the EQ stops. It is wrapped in foil, and ready for its journey into the IFO room. When we next have the BS chamber open, we should put the PRM in, and move the current PRM to its final place on the ITMY table as the SRM.